A water wave travels at a speed of 135 m/s, If the frequency of wave is 45 Hz what is the wavelength?
1
Expert's answer
2012-05-24T08:57:51-0400
A water wave travels at a speed of 135 m/s, If the frequency of wave
is 45 Hz
what is the wavelength?
V = 135 m/s.
F = 45 Hz = 45 waves/s.
L = V / F
= 135 / 45 = 3 m.